What is the [H+] of a solution with a pH of 9.7?
1
Expert's answer
2020-05-01T14:23:33-0400

pH=−log[H+]

9.7=−log[H+]

10−(9.7)=[H+]

[H+]=10−(9.7)=1.99×10−10 mol/L